# Volume effect on the extraction of sound velocity in high-energy nucleus-nucleus collisions

### v1

@@ Abstract
The determination of the speed of sound in quark-gluon plasma is a crucial aspect of understanding the properties of strongly interacting matter created in relativistic heavy-ion collisions. In this study, we investigate the impact of initial-state fluctuations on the extraction of the speed of sound in a quark-gluon plasma in the ultra-central collisions. By employing the TRENTo model for simulating initial conditions, we demonstrate that these fluctuations lead to sizable volume effect, which in turn corrects the measured values of the sound velocity. With respect to realistic conditions of high-energy heavy-ion experiments, we provide quantitative estimate of these corrections.
@@
$c_{s}^{2}\equiv\frac{dP}{de}\Big{|}_{X}\,.$ (1)
For instance, with respect to a quantum chromodynamic (QCD) system in thermal equilibrium, the solution of sound velocity from the lattice calculations of QCD implicitly assumes adiabatic condition Bazavov et al. (2014).
Speed of sound can be measured in experiments. Consider, for instance, ideally a thermal system in equilibrium is constrained in a box with fixed volume. As one warms up the system by generating entropy adiabatically, effectively the increase of temperature, $\Delta T$, and entropy $\Delta S$, can be used to estimate the sound velocity. By definition and the thermodynamic relations Van Hove (1982),
@@
Recently, in spirit of Eq. (2), the measurement of the speed of sound of a thermal QCD system has been proposed Gardim et al. (2020a) and carried out via the high-energy ultra-central collisions at the Large Hadron Collider (LHC) Hayrapetyan et al. (2024); Nielsen (2024). In these ultra-central collisions, system volume approaches maximum, while event-by-event quantum fluctuations from nucleon-nucleon collisions provide a means of entropy production adiabatically. In Eq. (2), the condition of thermal equilibrium is essential, although in realistic high-energy heavy-ion collisions, thermal equilibrium in the produced quark-gluon plasmas (QGP) is expected only locally, and even approximately, leading to a dissipative fluid description of the system expansion Shen and Yan (2020). Accordingly, the conservation of energy and entropy allows one to effectively convert the system into a homogenous QGP with volume $V_{\rm eff}$ and temperature $T_{\rm eff}$. Regarding final-state observables, hydrodynamic response relations are introduced to further relate the effective temperature $T_{\rm eff}$ to the mean value of the transverse momentum of the produced particles $\langle p_{T}\rangle$: $T_{\rm eff}\propto\langle p_{T}\rangle$, and the total entropy to the charged particle multiplicity: $S\propto dN_{\rm ch}/dy$. Eventually, the conjecture Gardim et al. (2020a) states that sound velocity in QGP can be measured in experiments from the ultra-central collisions by
@@
To avoid confusion, in Eq. (3) and the following, we denote the extracted sound velocity from the ratio of relative change of the effective temperature to that of the total entropy by $c_{s}^{0}$.
To validate this conjecture, several theoretical analyses have been performed Gardim et al. (2020b, 2024); Soares Rocha et al. (2024); Nijs and van der Schee (2024). In this work, assuming the well-established hydrodynamic response relations, we examine the extraction of sound velocity with respect to the fluctuating initial state of realistic heavy-ion collisions. On the first sight, it might seem trivial because the first equation in Eq. (3) follows simply the thermodynamic relation in Eq. (2). However, on an event-by-event basis, initial state density with fluctuations is neither homogeneous, nor sufficiently large that variation of the effective volume can be negligible. As a consequence, finite-volume effect arises which corrects the extracted sound velocity from Eq. (3), leading to,
@@
Here, the ratio between the relative change of volume and total entropy $\Delta\ln V_{\rm eff}/\Delta\ln S$, characterizes the response of the system volume as entropy is produced, which is not necessarily vanishing as one converts a fluctuating system to a homogeneous one. In fact, imposing the condition that $\Delta\ln V_{\rm eff}/\Delta\ln S=0$ induces violation of adiabiticity.
@@ II Effective modeling of initial state
III Extraction of sound velocity
Figure 1: (Color online) The initial-state parameter dependent relations between $T_{\rm eff}$ and $S$, and $V_{\rm eff}$ and $S$. Left panels ((a) and (b)): variation of width parameter $w$, with $p$ and $k$ fixed. Middle panels ((c) and (d)): variation of the fluctuation parameter $k$, with $p$ and $w$ fixed. Right panels ((e) and (f)): variation of the fluctuation parameter $p$, with $k$ and $w$ fixed.
The procedure of extracting sound velocity in the QGP medium starts with solving the effective temperature and effective volume Gardim et al. (2020a) from the equations,
@@ III Extraction of sound velocity
where the final-state total energy and entropy are estimated from particles emitted from the freeze-out hyper-surface. We employ a parameterized lattice equation of state (LEOS) that interpolates between the hadron resonance gas at low temperatures to the lattice results of QCD at high temperatures. With same parameters as those used in Bazavov et al. (2014), the effective parameterization of the LEOS has been widely applied in many hydrodynamic calculations.
With respect to the conservation of energy and entropy, the equations can be imposed as well if the total energy and entropy are given by the initial states,
@@
where the initial energy density is achieved by converting the entropy density using the LEOS, $e=e(s)$111 An alternative procedure is to relate the reduced thickness $T_{R}(x,y)$ to the initial energy density profile, then entropy density is obtained by a conversion via LEOS $s=s(e)$ Everett et al. (2021). . The initial proper time $\tau_{0}$ in hydrodynamic modeling is often considered with respect to a quick thermalization assumption, with $\tau_{0}\leq 1$ fm/c. In our analysis, the exact value of $\tau_{0}$ is irrelevant to the extraction of the speed of sound. The dependence on the space-time rapidity, $\eta_{s}={\rm arctanh}(z/t)$, reflects the breaking of boost invariance in high-energy collisions, which is necessary for the conservation of energy during medium expansion. In this work, we consider a factorized density and a two-parameter function $s(\tau_{0},\vec{x}_{\perp},\eta_{s})=s(\tau_{0},\vec{x}_{\perp})\times w(\eta_{s})$ for the longitudinal density profile Chatterjee and Bożek (2018),
@@
Ideally, with respect to the energy and entropy conservation, Eq. (8) and Eq. (9) are equivalent. However, it should be emphasized that, due to the finite dissipative effect in realistic QGP, from Eq. (9) the resulted relation between the effective temperature and entropy is not expected comparable to the relation between the mean transverse momentum and charged particle multiplicity in experiements222 Finite viscous effects lead to entropy production and a change in the mean transverse momentum. Accordingly, in the ultra-central collisions, the distributions of the mean transverse momentum and the charged particle multiplicity, are supposed to be distorted, in comparion to the distribution of $T_{\rm eff}$ and $S$ solved from Eq. (9). .
By imposing these identities in Eq. (8), as well in Eq. (9), the medium from realistic heavy-ion collisions is effectively converted to a homogeneous QGP in equilibrium, with volume $V_{\rm eff}$ and temperature $T_{\rm eff}$. In order to solve a sound velocity that is consistent with LEOS, the conversion must be adiabatic. Accordingly, volume between the two systems must differ. For instance, for inhomogeneous systems extra entropy is produced during the conversion, similar to the thermalization process. If there is no change of volume between the realistic QGP and the homogeneous QGP, without work there must be heat flow out of the system according to the first law of thermodynamics, thus adiabaticity condition is violated.
@@ III.1 Dependence on initial state model parameters
Figure 2: (Color online) Extracted speed of sound from directly the linear relation between $T_{\rm eff}$ and $S$, $(c_{s}^{0})^{2}$ (top panels), and the extraction with corrections from the volume effect, $c_{s}^{2}$ (bottom panels), as a function of the effective temperature. The orange line corresponds to the sound speed from LEOS. Dependence on the initial-state parameters, $w$, $k$, and $p$, are obvious in $(c_{s}^{0})^{2}$ in (a), (b), and (c), but significantly suppressed when volume effect is corrected, as in $c_{s}^{2}$ in (d), (e) and (f).
Without solving the entire medium evolution towards particlization, it is significantly simplified to extract the sound velocity from Eq. (9) on an event-by-event basis. In particular, the effect of initial state fluctuations on the extraction of the speed of sound can be studied. For each set of parameters in the TRENTo model, $(w,k,p)$, we select and simulate roughly five million events across the 0% to 5% centrality class of Pb-Pb collisions. We also vary the overall constant $C$ in Eq. (7), in order to scan a wider range of effective temperatures Gardim et al. (2024).
We first fix the fluctuation parameter $k$ and the generalized mean $p$ by their most probable values, and vary the effective nucleon width around $w=1.0$ fm, results are shown in the left panels of Fig. 1 ((a) and (d)). In analogy to the measurements of mean transverse momentum and multiplicity in experiments, we plot the effective temperature normalized by its average value in these ultra-central events, $T_{\rm eff}/T_{\rm eff}^{0}$, as a function of the total entropy scaled by its corresponding average, $S/S^{0}$. As shown in Fig. 1 (a), these are data points solved within approximately 40 bins in the 0%-5% centrality class. Statistical errors of these points are estimated with respect to their standard deviation. At $S>S^{0}$, the system size tends to saturate with a vanishing impact parameter, one observes a perfect linear relation between $T_{\rm eff}/T_{\rm eff}^{0}$ and $S/S^{0}$, with the slope corresponding to $c_{s}^{0}$ that can be extracted following Eq. (3). Note that, unlike the mean transverse momentum, from event-by-event simulations of initial states, there is no plateau and dip observed below $S^{0}$ in the effective temperature Nielsen (2024); Hayrapetyan et al. (2024); Nijs and van der Schee (2024). With the width growing from $w=0.5$ fm (blue points), to $w=1.0$ fm (brown points), to $w=1.5$ fm (green points), the resulted effective temperature decreases, but the slope increases.
@@ III.2 Volume effect
Figure 3: (Color online) Initial-state entropy density generated from the TRENTo model with the same seed generator to demonstrate the change of inhomogeneity with respect to the increase of $p$ and/or $w$.
One immediately deduces that $c_{s}^{0}$ barely relies on the fluctuation parameter $k$, as both $T_{\rm eff}$ and $S$ are not affected by the variation of $k$. However, the correlations of $T_{\rm eff}$ and $S$ to $w$ and $p$, result in strong initial-state parameter dependence of the extracted sound velocity $c_{s}^{0}$ from the event-by-event simulations. As shown in the top panels of Fig. 2, $c_{s}^{0}$, the slope extracted directed from the linearity between $T_{\rm eff}$ and $S$, captures to a large extent the physical speed of sound, as comparing to the input values from LEoS. However, the exacted values of $c_{s}^{2}$ can deviate from the physical speed of sound, depending on $p$ and $w$. For instance, at $T_{\rm eff}=225$ MeV, corresponding to the ultra-central PbPb collisions at $\sqrt{s_{NN}}=$ 5.02 TeV, small value of $p=-1$ gives rise to an underestimate of the speed of sound, while large $p$ leads to overestimation. In fact, one notices positive correlations between $(c_{s}^{0})^{2}$, and $p$ and $w$.
@@
Figure 4: (Color online) The ratio $\frac{\Delta\ln V}{\Delta\ln S}$ as a function of $N_{\rm ch}$ at different effective temperatures. Different points correspond to different initial-state parameters.
Figure 5: Inhomogeneous entropy production due to quantum fluctuations inside a system (shaded area) (I) or outside a system (shaded area) (II). When converting to a homogeneous thermal system with volume $V_{\rm eff}$ and entropy density $s$, a finite ratio $\frac{\Delta\ln V}{\Delta\ln S}$ is generated.
We now estimate the ratio $\Delta\ln V_{\rm eff}/\Delta\ln S$ for realistic high-energy heavy-ion experiments. Regarding the variation of $p$ or $k$ in the calculations, the effective temperatures $T_{\rm eff}$ can be determined with certain values. For all these fixed effective temperatures, the corresponding ratio $\Delta\ln V_{\rm eff}/\Delta\ln S$ exhibits linear correlation with the total entropy. Given the relation between the total entropy and the charged particle multiplicity, namely, $6.7\times dN_{\rm ch}/dy=dS/d\eta_{s}$ Hanus et al. (2019), these linear correlations are shown in Fig. 4. We find a formula that parameterizes all these lines,
$\frac{\Delta\ln V_{\rm eff}}{\Delta\ln S}=a-\exp\left[-\left(b_{0}+b_{1}T_{\rm eff}+b_{2}T_{\rm eff}^{2}\right)\right]N_{\rm ch}\,,$ (11)
with the constant coefficients: $a=1$, $b_{0}=-3.54$, $b_{1}=80.96$ GeV-1 and $b_{2}=-137.03$ GeV-2. Given Eq. (11), and the effective temperature from the measured mean transverse momentum, and the charged particle multiplicity in the ultra-central collisions, one is able to estimate the volume effect. The formula Eq. (11) should be valid for all the ultra-central nucleus-nucleus collisions carried out at the LHC and the top energy at the Relativistic Heavy-Ion Collider at the Brookhaven National Laboratory. In particular, for the measured effective temperature ($T_{\rm eff}=220$ MeV) and charged particle multiplicity ($dN_{\rm ch}/dy=1943\pm 54$ Adam et al. (2016)) regarding the ultra-central PbPb collisions at $\sqrt{s_{NN}}=$ 5.02 TeV at the LHC, Eq. (11) yields $\Delta\ln V_{\rm eff}/\Delta\ln S=0.065\pm 0.025$. Therefore, the measured value of the speed of sound Hayrapetyan et al. (2024); Nielsen (2024) is expected slightly smaller than the physical one.

### v3

# Volume effect on the extraction of speed of sound in high-energy nucleus-nucleus collisions

@@ Abstract
The determination of the speed of sound in quark-gluon plasma is a crucial aspect of understanding the properties of strongly interacting matter created in relativistic heavy-ion collisions. In this study, we investigate the impact of initial-state fluctuations on the extraction of the speed of sound in a quark-gluon plasma in the ultra-central collisions. By employing the TRENTo model for simulating initial conditions, we demonstrate that these fluctuations lead to sizable volume effect, which in turn corrects the measured values of the speed of sound. With respect to realistic conditions of high-energy heavy-ion experiments, we provide quantitative estimate of these corrections.
@@
$c_{s}^{2}\equiv\frac{dP}{de}\Big{|}_{X}\,,$ (1)
where $X$ denotes the conditions of, for instance, constant temperature, constant entropy, etc. In particular, with respect to a quantum chromodynamic (QCD) system in thermal equilibrium, the solution of the speed of sound from the lattice calculations of QCD implicitly assumes adiabatic condition Bazavov et al. (2014).
Speed of sound can be measured in experiments. Consider, for instance, ideally a thermal system in equilibrium is constrained in a box with fixed volume. As one warms up the system by generating entropy adiabatically, effectively the increase of temperature, $\Delta T$, and entropy $\Delta S$, can be used to estimate the speed of sound. By definition and the thermodynamic relations Van Hove (1982),
@@
Recently, in spirit of Eq. (2), the measurement of the speed of sound of a thermal QCD system has been proposed Gardim et al. (2020a) and carried out via the high-energy ultra-central collisions at the Large Hadron Collider (LHC) Hayrapetyan et al. (2024); Nielsen (2024). In these ultra-central collisions, system volume approaches maximum, while event-by-event quantum fluctuations from nucleon-nucleon collisions provide a means of entropy production adiabatically. In Eq. (2), the condition of thermal equilibrium is essential, although in realistic high-energy heavy-ion collisions, thermal equilibrium in the produced quark-gluon plasmas (QGP) is expected only locally, and even approximately, leading to a dissipative fluid description of the system expansion Shen and Yan (2020). Accordingly, the conservation of energy and entropy allows one to effectively convert the system into a homogenous QGP with volume $V_{\rm eff}$ and temperature $T_{\rm eff}$. Regarding final-state observables, hydrodynamic response relations are introduced to further relate the effective temperature $T_{\rm eff}$ to the mean value of the transverse momentum of the produced particles $\langle p_{T}\rangle$: $T_{\rm eff}\propto\langle p_{T}\rangle$, and the total entropy to the charged particle multiplicity: $S\propto dN_{\rm ch}/dy$. Eventually, the conjecture Gardim et al. (2020a) states that the speed of sound in QGP can be measured in experiments from the ultra-central collisions by
@@
To avoid confusion, in Eq. (3) and the following, we denote the extracted speed of sound from the ratio of relative change of the effective temperature to that of the total entropy by $c_{s}^{0}$.
To validate this conjecture, several theoretical analyses have been performed Gardim et al. (2020b, 2024); Soares Rocha et al. (2024); Nijs and van der Schee (2024). In this work, assuming the well-established hydrodynamic response relations, we examine the extraction of the speed of sound with respect to the fluctuating initial state of realistic heavy-ion collisions. On the first sight, it might seem trivial because the first equation in Eq. (3) follows simply the thermodynamic relation in Eq. (2). However, on an event-by-event basis, initial state density with fluctuations is neither homogeneous, nor sufficiently large that variation of the effective volume can be negligible. As a consequence, finite-volume effect arises which corrects the extracted speed of sound from Eq. (3), leading to,
@@
Here, the ratio between the relative change of volume and total entropy $\Delta\ln V_{\rm eff}/\Delta\ln S$, characterizes the response of the system volume as entropy is produced, which is not necessarily vanishing as one converts a fluctuating system to a homogeneous one. Note that, the condition $\Delta\ln V_{\rm eff}/\Delta\ln S=0$ implicitly states that there is no work being done to a system, as the system energy or entropy is varied during the measurement of the speed of sound, there must be heat transfer, hence adiabiticity is violated. More detailed discussions can be found in Appendix. A.
@@ II Effective modeling of initial state
III Extraction of the speed of sound
Figure 1: The linear correlations between initial- and final-state energy (left panel) and entropy (right panel).
Figure 2: The initial-state parameter dependent relations between $T_{\rm eff}$ and $S$, and $V_{\rm eff}$ and $S$. Left panels ((a) and (b)): variation of width parameter $w$, with $p$ and $k$ fixed. Middle panels ((c) and (d)): variation of the fluctuation parameter $k$, with $p$ and $w$ fixed. Right panels ((e) and (f)): variation of the fluctuation parameter $p$, with $k$ and $w$ fixed. Note that the points associated with temperatures below the cross-over temperature actually capture systems of hadron gas.
The procedure of extracting the speed of sound in the QGP medium starts with solving the effective temperature and effective volume Gardim et al. (2020a) from the equations,
@@ III Extraction of the speed of sound
where the final-state total energy and entropy are estimated from particles emitted from the freeze-out hyper-surface. A parameterized lattice equation of state (LEOS) that interpolates between the hadron resonance gas at low temperatures to the lattice results of QCD at high temperatures, relates the total energy and entropy to an effective temperature and an effective volume. With same parameters as those used in Bazavov et al. (2014), the effective parameterization of the LEOS has been widely applied in many hydrodynamic calculations.
The equations can be applied as well if the total energy and entropy are given by the initial states,
@@
where the initial energy density is achieved by converting the entropy density using the same LEOS, $e=e(s)$111 An alternative procedure is to relate the reduced thickness $T_{R}(x,y)$ to the initial energy density profile, then entropy density is obtained by a conversion via LEOS $s=s(e)$ Everett et al. (2021). . With respect to the total entropy, we are allowed to define centrality and to focus on the ultra-central events. Although the resulted extraction of the speed of sound is not identical between initial state and the final state, the two speeds are linked with each other as there exist linear correlations between initial state and final state, as shown in Fig. 1. Accordingly, it is still valuable to solve the speed of sound for the initial state, for which the analysis applies to a static thermal fireball subject to the QCD equation of state, in the presence of event-by-event fluctuations.
The initial proper time $\tau_{0}$ in hydrodynamic modeling is often considered with respect to a quick thermalization assumption, with $\tau_{0}\leq 1$ fm/c. In our analysis, the exact value of $\tau_{0}$ is irrelevant to the extraction of the speed of sound. The dependence on the space-time rapidity, $\eta_{s}={\rm arctanh}(z/t)$, reflects the breaking of boost invariance in high-energy collisions, which is necessary for the conservation of energy during medium expansion. In this work, we consider a factorized density and a two-parameter function $s(\tau_{0},\vec{x}_{\perp},\eta_{s})=s(\tau_{0},\vec{x}_{\perp})\times w(\eta_{s})$ for the longitudinal density profile Chatterjee and Bożek (2018),
@@
Ideally, with respect to the energy and entropy conservation, Eq. (8) and Eq. (9) are equivalent. However, it should be emphasized that, due to the finite dissipative effect in realistic QGP, from Eq. (9) the resulted relation between the effective temperature and entropy is not expected comparable to the relation between the mean transverse momentum and charged particle multiplicity in experiements [^2] Finite viscous effects lead to entropy production and a change in the mean transverse momentum. Accordingly, in the ultra-central collisions, the distributions of the mean transverse momentum and the charged particle multiplicity, are supposed to be distorted, in comparion to the distribution of $T_{\rm eff}$ and $S$ solved from Eq. (9). .
By imposing these identities in Eq. (8), as well in Eq. (9), the medium from realistic heavy-ion collisions is effectively converted to a homogeneous QGP in equilibrium, with volume $V_{\rm eff}$ and temperature $T_{\rm eff}$. In order to solve a speed of sound that is consistent with LEOS, the conversion must be adiabatic. Accordingly, volume between the two systems must differ. For instance, for inhomogeneous systems extra entropy is produced during the conversion, similar to the thermalization process. If there is no change of volume between the realistic QGP and the homogeneous QGP, without work there must be heat flow out of the system according to the first law of thermodynamics, thus adiabaticity condition is violated.
@@ III.1 Dependence on initial state model parameters
Figure 3: Extracted speed of sound from directly the linear relation between $T_{\rm eff}$ and $S$, $(c_{s}^{0})^{2}$ (top panels), and the extraction with corrections from the volume effect, $c_{s}^{2}$ (bottom panels), as a function of the effective temperature. The orange line corresponds to the sound speed from LEOS. Dependence on the initial-state parameters, $w$, $k$, and $p$, are obvious in $(c_{s}^{0})^{2}$ in (a), (b), and (c), but significantly suppressed when volume effect is corrected, as in $c_{s}^{2}$ in (d), (e) and (f).
Without solving the entire medium evolution towards particlization, it is significantly simplified to extract the speed of sound from Eq. (9) on an event-by-event basis. In particular, the effect of initial state fluctuations on the extraction of the speed of sound can be studied. For each set of parameters in the TRENTo model, $(w,k,p)$, we select and simulate roughly five million events across the 0% to 5% centrality class of Pb-Pb collisions. We also vary the overall constant $C$ in Eq. (7), in order to scan a wider range of effective temperatures Gardim et al. (2024).
We first fix the fluctuation parameter $k$ and the generalized mean $p$ by their most probable values, and vary the effective nucleon width around $w=1.0$ fm, results are shown in the left panels of Fig. 2 ((a) and (d)). In analogy to the measurements of mean transverse momentum and multiplicity in experiments, we plot the effective temperature normalized by its average value in these ultra-central events, $T_{\rm eff}/T_{\rm eff}^{0}$, as a function of the total entropy scaled by its corresponding average, $S/S^{0}$. As shown in Fig. 2 (a), these are data points solved within approximately 40 bins in the 0%-5% centrality class. Statistical errors of these points are estimated with respect to their standard deviation. At $S>S^{0}$, the system size tends to saturate with a vanishing impact parameter, one observes a perfect linear relation between $T_{\rm eff}/T_{\rm eff}^{0}$ and $S/S^{0}$, with the slope corresponding to $c_{s}^{0}$ that can be extracted following Eq. (3). Note that, unlike the mean transverse momentum, from event-by-event simulations of initial states, there is no plateau and dip observed below $S^{0}$ in the effective temperature Nielsen (2024); Hayrapetyan et al. (2024); Nijs and van der Schee (2024). The plateau and dip structure which reflects the residual dependence on the system size of the mean transverse momentum, disappears as the system size saturates when centrality approach a limit, e.g., $S>S^{0}$. Similar effect can be observed in $T_{\rm eff}$ that violation of the linearity exhibits in the smaller entropy region. Nonetheless, the extraction of the speed of sound does not rely on these structures. With the width growing from $w=0.5$ fm (blue points), to $w=1.0$ fm (brown points), to $w=1.5$ fm (green points), the resulted effective temperature decreases, but the slope increases.
@@ III.2 Volume effect
Figure 4: Initial-state entropy density generated from the TRENTo model with the same seed generator to demonstrate the change of inhomogeneity with respect to the increase of $p$ and/or $w$.
One immediately deduces that $c_{s}^{0}$ barely relies on the fluctuation parameter $k$, as both $T_{\rm eff}$ and $S$ are not affected by the variation of $k$. However, the correlations of $T_{\rm eff}$ and $S$ to $w$ and $p$, result in strong initial-state parameter dependence of the extracted speed of sound $c_{s}^{0}$ from the event-by-event simulations. As shown in the top panels of Fig. 3, $c_{s}^{0}$, the slope extracted directed from the linearity between $T_{\rm eff}$ and $S$, captures to a large extent the physical speed of sound, as comparing to the input values from LEoS. However, the exacted values of $c_{s}^{2}$ can deviate from the physical speed of sound, depending on $p$ and $w$. For instance, at $T_{\rm eff}=225$ MeV, corresponding to the ultra-central PbPb collisions at $\sqrt{s_{NN}}=$ 5.02 TeV, small value of $p=-1$ gives rise to an underestimate of the speed of sound, while large $p$ leads to overestimation. In fact, one notices positive correlations between $(c_{s}^{0})^{2}$, and $p$ and $w$.
@@
Figure 5: Inhomogeneous entropy production due to quantum fluctuations inside a system (shaded area) (I) or outside a system (shaded area) (II). When converting to a homogeneous thermal system with volume $V_{\rm eff}$ and entropy density $s$, a finite ratio $\alpha=\frac{\Delta\ln V}{\Delta\ln S}$ is generated.